Ask
your own question, for FREE!
Mathematics
11 Online
ok so You roll one number cube twice. Find each probability. 1: P(1 and odd)
Still Need Help?
Join the QuestionCove community and study together with friends!
How do i do this?
First find the probability of P(1), which is 1/6 Then find P(odd), which is 3/6 = 1/2
But how did u get 3/6?
Multiply the two probabilities together to find the probability they will happen in sequence. 1/6 * 1/2 = 1/12
Wait im Confused what am i Suppose to Rigth?
Still Need Help?
Join the QuestionCove community and study together with friends!
I found 3/6 because there are 3 odd numbers out of 6 total
and 1/6 because there is one 1 and 6 total numbers on a die
the final answer is 1/12
so i write to answer it P(1/6x1/2=1/12?
P(1) = 1/6 P(odd) = 1/2 P(1)*P(odd) = 1/12
Still Need Help?
Join the QuestionCove community and study together with friends!
Can uAlso Help me with this P(3 and even is the even 2?
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
11 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal